Q. A proton and an alpha particle moving with same velocity in a uniform magnetic field which is acting normally to their plane of motion. The ratio of the radii of the circular paths described by the proton and alpha particle respectively is what?

Grade:12th pass

1 Answers

hero
7 years ago
Magnetic force=centripetal force.qvB=mv^2÷rOr r=mv÷qBv-same. B-same therefore ;r(p)/r(a)=m(p)×q(a)/m(a)×q(p) On solving ratio = 1:2